System of Linear Equations:
Solve the system using the addition method.
5x=6y+40
2y=8-3x

Answer by checkley77(12844)

5x=6y+40
2y=8-3x MULTIPLY THIS EQUATION BY -3 & ADD.
-6y=-24+9X
6y=-40+5x
-------------------
0y=-64+14x
14x=64
x=64/14
x=32/7 answer.
5*32/7=6y+40
160/7=6y+40
6y=160/7-40
6y=(160-40*7)/7
6y=(160-280)/7
6y=-120/7
y=-120/7*1/6
y=-120/42
y=-20/7 answer.